OpenStudy (brinazarski):

The perimeter of a rectangle is 52 inches, and its area is 168 in2. Find its length and width.

OpenStudy (anonymous):

 so perimeter is 2x+2y which equals 52 inches 2x+2y=52 area is xy and that equals 168 xy=168 now you have two unknowns, two equations, any idea what to do from here?

OpenStudy (brinazarski):

2x+(168/x)=52?

OpenStudy (anonymous):

looks good so far, keep going

OpenStudy (brinazarski):

2(x^2-26x+84)=0?

OpenStudy (brinazarski):

I think that's where I got lost last time I tried to solve this

OpenStudy (brinazarski):

I retried it and got x = sqrt(85) +13... but I'm doing this online and this problem says it's going to be rational number

OpenStudy (anonymous):

Just a second gonna do it on paper

OpenStudy (brinazarski):

Okay

OpenStudy (anonymous):

ohhh you left off the 2 thats why i got messed up. I shoulda looked at original problem. anyways, 2x+2(168/x)=52 2x+336/x=52 (2x^2/x)+(336/x)=52 (2x^2+336)/x=52 2x^2+336=52x 2x^2-52x+336=0 ok, see how i got this far?

OpenStudy (brinazarski):

Yup!

OpenStudy (anonymous):

factor out a two 2(x^2-26x+168)=0 now factor whichever way you like best

OpenStudy (brinazarski):

I got the answer from there :) Thank you so much!

OpenStudy (anonymous):

You're welcome